Blossom Company makes three models of tasers. Information on the three products is given below.

Tingler Shocker Stunner
Sales $304,000 $496,000 $200,000
Variable expenses 154,500 190,800 143,800
Contribution margin 149,500 305,200 56,200
Fixed expenses 120,200 228,900 94,400
Net income $29,300 $76,300 $(38,200)


Fixed expenses consist of $300,000 of common costs allocated to the three products based on relative sales, as well as direct fixed expenses unique to each model of $29,000 (Tingler), $80,100 (Shocker), and $34,400 (Stunner). The common costs will be incurred regardless of how many models are produced. The direct fixed expenses would be eliminated if that model is phased out.

James Watt, an executive with the company, feels the Stunner line should be discontinued to increase the company’s net income.

(a)

Compute current net income for Blossom Company.

Net income $

Entry field with correct answer


(b)

Compute net income by product line and in total for Blossom Company if the company discontinues the Stunner product line. (Hint: Allocate the $300,000 common costs to the two remaining product lines based on their relative sales.)

Tingler Net Income $

Entry field with incorrect answer now contains modified data

Shocker Net Income $

Entry field with incorrect answer now contains modified data

Total Net Income $

Entry field with incorrect answer now contains modified data


(c)

Should Blossom eliminate the Stunner product line?

Entry field with correct answer YesNo



Why or why not?

Net income would

Entry field with correct answer decreaseincrease

from $

Entry field with correct answer

to $

Entry field with incorrect answer now contains modified data

.
Answer #1
a
Net income 67400
b
Tingler Net Income 6500
Shocker Net Income 39100
Total Net Income 45600
c
NO, do not eliminate the Stunner product line
Net income would decrease from $67400 to $45600
Workings:
a
Tingler net income 29300
Shocker net income 76300
Stunner net income -38200
Net income 67400
b
Tingler Shocker Total
Sales 304000 496000 800000
Variable expenses 154500 190800 345300
Contribution margin 149500 305200 454700
Fixed expenses 143000 266100 409100
Net income 6500 39100 45600
Fixed expenses :
Tingler=29000+(300000/800000*304000) = $143000
Shocker=80100+(300000/800000*496000) = $266100